Grade 11General Physics

Two particles P and Q are moving with velocities of (unit vector i+ unit vector j) and (unit vector -i + unit vector 2j) respectively. At time t=0, P is at Origin and Q is at a point with position vector (unit vector 2i+ unit vector j). Then the shortest distance between P and Q is?

Diya Sharma
10 Years agoGrade 11

1 Answer

Akshay
10 Years ago
Relative velocity of Q wrt P = (-2i + j), Relative position of Q wrt P = (2i + j). Path of Q in frame of P will touch (0,2) point and (2,1) point
So Q will move along a line 2y+x=4 wrt P.
Shortest distance between P and Q = 4/rt(4+1) = 4/rt(5).........which is perpendicular distance between (0,0)  and 2y+x=4.